### v1

# Entanglement between dependent degrees of freedom: Quasi-particle correlations

Abstract
Common notions of entanglement are based on well-separated subsystems. However, obtaining such independent degrees of freedom is not always possible because of physical constraints. In this work, we explore the notion of entanglement in the context of dependent degrees of freedom. As a physically relevant application, we specifically study quantum correlation features for quasi-particle descriptions. Those are paramount for interacting light-matter systems, utilizing excitations of fermion-boson hybrid modes. By comparing independent and dependent degrees of freedom, we uncover that certain states are non-entangled although they would be entangled when only focusing on the common, independent description, and vice versa. Therefore, new insight is provided into the resourcefulness of quantum correlations within the rarely discussed context of dependent degrees of freedom for light-matter links in quantum information applications.
I Introduction
Features that surpass classical limitations are at the heart of quantum science and quantum technologies [1, 2]. Arguably the most versatile quantum phenomenon is entanglement [3], being crucial for many applications in quantum communication and quantum computation [4], such as teleportation [5] and dense coding [6]. While determining whether a state is entangled is known to be a computationally hard problem [7], certain states are easily recognizable as entangled, e.g., W states, GHZ states, and NOON states [8, 9, 10].
Entanglement in hybrid systems [11, 12] is crucial for quantum links between distinct physical platforms, e.g., using matter properties for processing and light for communication [13, 12]. The quasi-particle description is a convenient way to describe such hybrid systems, including the assessment of quantum resources stemming from light-matter interactions [15, 14]. Interfacing bosonic and fermionic parts in quantum links, quasi-particles thus constitute joint excitations of quantized bosonic and fermionic modes resulting in mixed quantum statistics for the quasi-particles [16].
Quantum properties of quasi-particles, including entanglement, have been studied [17, 15, 18, 19, 20, 14]. Still, the versatile nature of entanglement means that different forms of entanglement exist [21, 22, 23, 24]. For example, it can occur independently for particles and fields to the point that both forms can be classified as two distinct phenomena [25]. In addition, entanglement can be independent of the decomposition of the full system into quantized basis modes, allowing entanglement to persist regardless the chosen mode basis [26]. This and other findings raise the question of whether the common analysis of entanglement applies to the quasi-particle pictures, too.
In optics, for example, commonly studied degrees of freedom include, but are not limited to orbital angular momentum, time-bin modes, spatial modes, and polarization [27]. In all these instances, entanglement is formulated in terms of independent degrees of freedom of purely bosonic excitations [28, 29, 30, 31, 32]. However, degrees of freedom can be dependent because of physical constraints, which is not commonly discussed and the focus of this work.
In this work, we overcome the aforementioned limitation by characterizing entangled and non-entangled states between dependent degrees of freedom. This is mainly done for light-matter interfaces and their quasi-particle descriptions via a mode transformation to quasi-particle excitations arising from diagonalizing a Hamiltonian in terms of quantized fields. We show instances of both entangled and non-entangled states for dependent degrees of freedom whose independent counterparts behave oppositely, showing the distinctiveness of quantum correlations in such scenarios. Also, we compare the fermion-boson quasi-particle modes with analogous models consisting purely of bosonic and fermionic subsystems. Finally, a generalization to other scenarios and instances of multipartite entanglement is outlined.
The remainder of the work is structured as follows. In Sec. II, we introduce the specific system under study. Entanglement for quasi-particle descriptions is established and investigated in Sec. III. A comparison with independent degrees of freedom, i.e., purely bosonic and purely fermionic scenarios, is discussed in Sec. IV. In Sec. V, we formalize and generalize our approach. We conclude in Sec. VI.
II Light-matter interface
We begin by introducing the specific system under study. This particular system functions as a well-studied, proof-of-concept example for analyzing entanglement of independent and dependent degrees of freedom throughout this work.
II.1 Quantized eigenmodes of a Hamiltonian
Consider a composite fermion-boson system in second quantization with independent field operators $\hat{f}$ and $\hat{b}$, respectively, obeying $\{\hat{f},\hat{f}\}=0=[\hat{b},\hat{b}]$ and $\{\hat{f},\hat{f}^{\dagger}\}=\hat{1}=[\hat{b},\hat{b}^{\dagger}]$, where $\hat{1}$ denotes the identity. The system’s Hamiltonian under study consists of free-field contributions and couplings, analogous to the Jaynes–Cummings model [33], reading
$\frac{\hat{H}}{\hbar}=\omega_{F}\hat{f}^{\dagger}\hat{f}+\omega_{B}\hat{b}^{\dagger}\hat{b}+\frac{\kappa}{2}\hat{f}^{\dagger}\hat{b}+\frac{\kappa^{\ast}}{2}\hat{b}^{\dagger}\hat{f},$ (1)
with a complex coupling parameter $\kappa\neq 0$. It is convenient to introduce the abbreviations
$\Omega=\omega_{F}+\omega_{B}\quad\text{and}\quad\delta=\omega_{F}-\omega_{B},$ (2)
where $\delta$ is dubbed the detuning. Using an $\mathrm{SU}(2)$ map, we can find the eigenmodes of the Hamiltonian, here named $\hat{p}_{\pm}$, as
$\begin{bmatrix}\hat{p}_{+}\\ \hat{p}_{-}\end{bmatrix}=\begin{bmatrix}\varphi&\beta e^{i\vartheta}\\ -\beta e^{-i\vartheta}&\varphi\end{bmatrix}\begin{bmatrix}\hat{f}\\ \hat{b}\end{bmatrix},$ (3)
where
$\varphi=\sqrt{\frac{\Delta+\delta}{2\Delta}},\quad\beta=\sqrt{\frac{\Delta-\delta}{2\Delta}},\quad\text{and}\quad\frac{\kappa}{|\kappa|}=e^{i\vartheta},$ (4)
with $\Delta=\sqrt{\delta^{2}+|\kappa|^{2}}$ and $\varphi^{2}+\beta^{2}=1$. Excitations of these eigenmodes define quasi-particles, which we later discuss in detail. See also Ref. [34] in this context. Furthermore, we can recast the Hamiltonian in Eq. (1) in a noninteracting form as
$\hat{H}=E_{+}\hat{p}_{+}^{\dagger}\hat{p}_{+}+E_{-}\hat{p}_{-}^{\dagger}\hat{p}_{-},\quad\text{with}\quad E_{\pm}=\frac{\hbar}{2}\left(\Omega\pm\Delta\right)$ (5)
and without cross-terms between the quasi-particle modes; see also Fig. 1 for the energies pertaining to the eigenmodes.
Figure 1: Energy bands $E_{\pm}$ from diagonalization of the Hamiltonian via the introduction of quasi-particle modes $\hat{p}_{\pm}$ in Eq. (5) as a function of the ratio of coupling strength $|\kappa|$ and detuning $\delta$.
II.2 Quantum states of the system
Via the vacuum state $|\mathrm{vac}\rangle$, the Fock basis in the boson-fermion description consists of the elements
$|0,n\rangle_{\mathrm{FB}}=\frac{\hat{b}^{{\dagger}n}}{\sqrt{n!}}|\mathrm{vac}\rangle\quad\text{and}\quad|1,n\rangle_{\mathrm{FB}}=\hat{f}^{\dagger}\frac{\hat{b}^{{\dagger}n}}{\sqrt{n!}}|\mathrm{vac}\rangle,$ (6)
for $n\in\mathbb{N}$ and utilizing $\hat{f}^{{\dagger}m}=0$ for all $m>1$. In addition to the vacuum state, all eigenstates (also known as dressed states [35]) of the Hamiltonian can be expanded as
$|\phi_{+,N}\rangle=$ ${}\beta_{N}|0,N\rangle_{\mathrm{FB}}+\varphi_{N}e^{i\vartheta}|1,N-1\rangle_{\mathrm{FB}}$ (7)
$\text{and}\quad|\phi_{-,N}\rangle=$ ${}\varphi_{N}|0,N\rangle_{\mathrm{FB}}-\beta_{N}e^{i\vartheta}|1,N-1\rangle_{\mathrm{FB}},$
with Rabi frequencies $\Delta_{N}=\sqrt{\delta^{2}+N|\kappa|^{2}}$ and parameters
$\varphi_{N}=\sqrt{\frac{\Delta_{N}+\delta}{2\Delta_{N}}}\quad\text{and}\quad\beta_{N}=\sqrt{\frac{\Delta_{N}-\delta}{2\Delta_{N}}}.$ (8)
Note that, for the single-particle excitation, $N=1$, we have $\Delta_{1}=\Delta$, $\varphi_{1}=\varphi$, and $\beta_{1}=\beta$, according to our definitions in Eq. (4). Moreover, the energies pertaining to the eigenstates in Eq. (7) are
$E_{\pm,N}=\frac{\hbar}{2}\left(N\Omega-N\delta+\delta\pm\Delta_{N}\right).$ (9)
The notion of entanglement in the original $|m,n\rangle_{\mathrm{FB}}$ is given in terms of independent degrees of freedom. However, we are going to explore entanglement with respect to the eigenmodes of the Hamiltonian, in which we separate quasiparticles excited through $\hat{p}_{+}$ from the excitations of $\hat{p}_{-}$. This is going to lead us to the notion of entanglement between dependent degrees of freedom.
III Quasi-particle entanglement
In this section, we establish and analyze the notion of entanglement for dependent degrees of freedom in the context of the model under study. We show how linear dependence plays a central role in this concept. Also, we explore how the different notions of entanglement for states of a given total particle number behave differently for independent and dependent degrees of freedom.
III.1 Product states in fermion-boson hybrid descriptions
For defining entanglement of dependent degrees of freedom, we firstly need to recapitulate non-entangled states for independent degrees of freedom. Specifically, product states of independent fermions and bosons take the form
$|x,y\rangle_{\mathrm{FB}}=$ ${}\left(\sum_{m=0}^{1}x_{m}|m\rangle_{\mathrm{F}}\right)\otimes\left(\sum_{n=0}^{\infty}y_{m}|m\rangle_{\mathrm{B}}\right)$ (10)
$=$ ${}\sum_{m=0}^{1}\sum_{n=0}^{\infty}\tilde{x}_{m}\tilde{y}_{n}\hat{f}^{{\dagger}m}\hat{b}^{{\dagger}n}|\mathrm{vac}\rangle,$
where $\tilde{x}_{m}=x_{m}$ and $\tilde{y}_{n}=y_{n}/\sqrt{n!}$ relate expansion coefficients in Fock basis and excitations of quantized fields. As examples, the basis states in Eq. (6) are classical states with respect to the fermion-boson separation. A state $|\psi\rangle$ that cannot be expressed as defined in Eq. (10) is, by definition, entangled.
Note that, throughout most parts of this work, we focus on pure states. However, it is worth mentioning that classically correlated, i.e., separable states [36], can be introduced by using an ensemble of pure product states in form of a density matrix. Still, all concepts discussed here can be extended to mixed, inseparable states, as discussed in Sec. V.
III.2 Field-product states for dependent degrees of freedom
Generalizing the second line in Eq. (10) renders it possible to define non-entangled states for the quasi-particle modes $\hat{p}_{\pm}$, Eq. (3). That is, a state is said to be non-entangled in this $\pm$ description if it can be expressed in a factorized manner as follows:
$|x,y\rangle_{\pm}=\left(\sum_{m\in\mathbb{N}}\tilde{x}_{m}\hat{p}_{+}^{{\dagger}m}\right)\left(\sum_{n\in\mathbb{N}}\tilde{y}_{n}\hat{p}_{-}^{{\dagger}n}\right)|\mathrm{vac}\rangle,$ (11)
with complex coefficients $x_{m}$ and $y_{n}$ such that the state is normalized, ${}_{\pm}\langle x,y|x,y\rangle_{\pm}=1$. Again, a convex mixture in terms of a density matrix extends this notion to separable mixed states in the $\pm$ separation.
Similarly to the previous case, and according to the definition in Eq. (11), the following $N$-particle states, where $N=m+n$, are product states:
${}\hat{p}_{+}^{{\dagger}m}\hat{p}_{-}^{{\dagger}n}|\mathrm{vac}\rangle$ (12)
$=$ ${}e^{-im\vartheta}\left(\beta^{m}\hat{b}^{{\dagger}m}+m\varphi\beta^{m-1}e^{i\vartheta}\hat{f}^{\dagger}\hat{b}^{{\dagger}[m-1]}\right)$
×${}\times\left(\varphi^{n}\hat{b}^{n}-n\beta\varphi^{n-1}e^{i\vartheta}\hat{f}^{\dagger}\hat{b}^{{\dagger}[n-1]}\right)|\mathrm{vac}\rangle$
$=$ $e^{-im\vartheta}\beta^{m}\varphi^{n}\sqrt{N!}$
×${}\times\left(|0,N\rangle_{\mathrm{FB}}+\frac{m\varphi^{2}-n\beta^{2}}{\beta\varphi\sqrt{N}}e^{i\vartheta}|1,N-1\rangle_{\mathrm{FB}}\right),$
being also expanded in the $\mathrm{FB}$ basis for comparison. Including a proper normalization and ignoring global phases, this yields the $\pm$-product states
$|m,n\rangle_{\pm}=\frac{|0,N\rangle_{\mathrm{FB}}+\frac{m\varphi^{2}-n\beta^{2}}{\beta\varphi\sqrt{N}}e^{i\vartheta}|1,N-1\rangle_{\mathrm{FB}}}{\sqrt{1+\frac{(m\varphi^{2}-n\beta^{2})^{2}}{\beta^{2}\varphi^{2}N}}}.$ (13)
Note that, unsurprisingly, such states are generally product states with respect to the quasi-particle modes $\hat{p}_{\pm}$ but entangled with respect to the fermion-boson separation $\mathrm{FB}$ since a global mode transformation is used in Eq. (3). Also, we can choose $m,n\in\mathbb{N}$ since neither $\hat{p}_{+}^{{\dagger}m}$ nor $\hat{p}_{-}^{{\dagger}n}$ become zero for a finite integer, contrasting pure fermions.
III.3 Linear dependence and disentangled NOON states
It is straightforward to observe that, for any two pairs of integers $(m,n)$ and $(m^{\prime},n^{\prime})$, which obey $(m\varphi^{2}-n\beta^{2})/\sqrt{m+n}\neq(m^{\prime}\varphi^{2}-n^{\prime}\beta^{2})/\sqrt{m^{\prime}+n^{\prime}}$, the states $|m,n\rangle_{\pm}$ and $|m^{\prime},n^{\prime}\rangle_{\pm}$ are linearly independent. However, for a fixed $N>0$, there are only two linearly independent elements in total since the $\mathrm{FB}$ expansion is based on two basis elements only; see Eq. (13). Therefore, the identities
$|1,N-1\rangle_{\mathrm{FB}}=$ ${}\frac{e^{-i\vartheta}\varphi\beta}{\sqrt{N}}\left(\sqrt{1+N\frac{\varphi^{2}}{\beta^{2}}}|N,0\rangle_{\pm}\right.$ (14)
${}\left.-\sqrt{1+N\frac{\beta^{2}}{\varphi^{2}}}|0,N\rangle_{\pm}\right)$
and
$|0,N\rangle_{\mathrm{FB}}=$ ${}\varphi^{2}\sqrt{1+N\frac{\beta^{2}}{\varphi^{2}}}|0,N\rangle_{\pm}$ (15)
${}+\beta^{2}\sqrt{1+N\frac{\varphi^{2}}{\beta^{2}}}|N,0\rangle_{\pm}$
imply for all $m=N-n$ and $n\notin\{0,N\}$ that we find
$|m,n\rangle_{\pm}=$ ${}\frac{m}{N}\sqrt{\frac{N\varphi^{2}\beta^{2}+N^{2}\varphi^{4}}{N\varphi^{2}\beta^{2}+\left(m\varphi^{2}-n\beta^{2}\right)^{2}}}|N,0\rangle_{\pm}$ (16)
${}+\frac{n}{N}\sqrt{\frac{N\varphi^{2}\beta^{2}+N^{2}\beta^{4}}{N\varphi^{2}\beta^{2}+\left(m\varphi^{2}-n\beta^{2}\right)^{2}}}|0,N\rangle_{\pm}.$
For $N>2$, this shows that all but two $\pm$-product states can be expressed via superpositions of only two elements.
As a consequence of the above linear combination, so-called NOON states [9, 10],
$|N00N\rangle_{\pm}=\psi_{0}|N,0\rangle_{\pm}+\psi_{N}|0,N\rangle_{\pm},$ (17)
with $\psi_{0}\psi_{N}\neq 0$, are not necessarily entangled for $N>2$. Recall that, by contrast, all NOON states are entangled for independent degrees of freedom. Here, whenever the coefficients $\psi_{0}$ and $\psi_{N}$ take the form of Eq. (16), such NOON states clearly admit a product form $|m,n\rangle_{\pm}$ for the dependent degrees of freedom $\pm$. Again, we emphasize that this disentanglement of quasi-particle NOON states is distinct from common discussions of entanglement. See Fig. 2
Figure 2: NOON-type states from Eq. (17) on the circle in an FB coordinate system, for $N=5$ quasi-particles, where $\varphi=\beta=1/\sqrt{2}$ and $\vartheta=0$. All non-entangled states in the $\pm$ description, Eq. (16), are shown as bullet points, including a $\pm$ coordinate system of the dependent degrees of freedom, which is indicated through dashed lines. Note that $N+1=6$ states, plus their reflections about the origin, are shown in the real plane, ignoring complex phases of $\psi_{0}$ and $\psi_{N}$ for simplicity. For independent degrees of freedom, all but the states $|N,0\rangle_{\pm}$ and $|0,N\rangle_{\pm}$ (and their reflections $-|N,0\rangle_{\pm}$ and $-|0,N\rangle_{\pm}$) alone would be entangled.
III.4 Quantum-correlated eigenstates of non-interacting Hamiltonians
Next, consider the non-interacting case, where $\kappa=0$ for the Hamiltonian in Eq. (1), and the independent degrees of freedom $\mathrm{FB}$. This yields a Hamiltonian which decomposes into the single-mode parts $\hbar\omega_{F}\hat{f}^{\dagger}\hat{f}$ and $\hbar\omega_{B}\hat{b}^{\dagger}\hat{b}$, as well as vanishing interactions, $\hat{H}_{\mathrm{int}}=0$. In that case, the eigenstates are necessarily products, $|m,n\rangle_{\mathrm{FB}}$ [Eq. (6)], for the non-degenerate case $\delta\neq 0$. This is generally true for independent degrees of freedom and $\hat{H}=\hat{H}_{1}\otimes\hat{1}_{2}+\hat{1}_{1}\otimes\hat{H}_{2}$, with subsystem identities $\hat{1}_{j}$ and local Hamiltonians $\hat{H}_{j}$ for $j\in\{1,2\}$. Further, the quasi-particle modes $\hat{p}_{\pm}$ were introduced to achieve a non-interacting representation of the Hamiltonian for $\kappa\neq 0$, Eq. (5). So the expectation derived from independent degrees of freedom would be that the $N$-particle eigenstates in Eq. (7) are products with respect to the $\pm$ separation.
By equating coefficients of the eigenstates of $\hat{H}$ and the $N$-particle product states in Eq. (13), we find the following two possible conditions for factorization:
$\frac{\varphi_{N}}{\beta_{N}}=\frac{m\varphi^{2}-n\beta^{2}}{\varphi\beta\sqrt{N}}\quad\text{or}\quad-\frac{\beta_{N}}{\varphi_{N}}=\frac{m\varphi^{2}-n\beta^{2}}{\varphi\beta\sqrt{N}},$ (18)
while recalling $\varphi_{1}=\varphi$, $\beta_{1}=\beta$, $N=m+n$, and Eq. (8). For $N>1$, those conditions are generally not met, therefore demonstrating that the eigenstates of an interaction-free Hamiltonian in dependent degrees of freedom can, in fact, be entangled. For example, the case $\delta=0$ applied to the conditions in Eq. (18) simplify them to
$m=\frac{N\pm\sqrt{N}}{2}\quad\text{and}\quad n=\frac{N\mp\sqrt{N}}{2},$ (19)
which has integer solutions if and only if the total particle number $N$ is a perfect square. Conversely, $\sqrt{N}\notin\mathbb{N}$ leads to entangled $N$-quasi-particle eigenstates, contrary to the separation of eigenstates for independent degrees of freedom without coupling terms in the Hamiltonian.
III.5 Preliminary discussion
We established the notion of classically correlated states for dependent degrees of freedom, being based on products states in Eq. (11). Formally, those states admit the same structure as non-entangled states for independent degrees of freedom, Eq. (10). However, we proved by examples two key distinguishing features of the different kinds of degrees of freedom.
In the first example, we showed linear dependence in the dependent case. This led to the observation of disentangled NOON states, which are always entangled for independent degrees of freedom. The second example pertained to the product nature of eigenstates of non-interacting Hamiltonians, which are products for independent degrees of freedom. Again, a departure from this behavior was observed for dependent degrees of freedom.
Therefore, quantum correlations of dependent degrees of freedom defy common assumptions about quantum correlations in a given system. This is interesting, e.g., in the context of light-matter interfaces for quantum information applications which require a firm understanding of quantum correlations for applications in various protocols.
IV Comparing different quantum statistics
Thus far, we compared the separation with respect to independent degrees of freedom, consisting of one fermionic and one bosonic field (i.e., $\mathrm{FB}$ separation), and the quasi-particle description via hybrid eigenmodes (i.e., a separation with respect to the $+$ and -- modes), obeying neither a fermionic nor a bosonic algebra. In this section, we directly compare the previous cases with composite systems consisting purely of fermionic and bosonic quantized fields.
IV.1 Fermions, bosons, and combinations thereof
For the sake of simplicity, we assume a vanishing detuning, $\delta=0$ (likewise, $\omega_{F}=\omega=\omega_{B}$), and a zero phase for the coupling parameter, $\kappa/|\kappa|=e^{i\theta}=1$, throughout this section. For example, this yields balanced superposition [see Eq. (3)] for the quasi-particle modes
$\hat{p}_{\pm}=\frac{\hat{f}\pm\hat{b}}{\sqrt{2}}.$ (20)
The factorizable $N$-quasi-particle states from Eq. (13) simplify to
$|m,n\rangle_{\pm}=\frac{|0,N\rangle_{\mathrm{FB}}+\frac{m-n}{\sqrt{N}}|1,N-1\rangle_{\mathrm{FB}}}{\sqrt{1+\frac{(m-n)^{2}}{N}}},$ (21)
where $N=m+n$. Note that the vacuum state is trivially the same for all representations, $|\mathrm{vac}\rangle=|0,0\rangle_{\mathrm{FB}}=|0,0\rangle_{\pm}$.
Analogously to the previous scenario, Eq. (1), we can consider Hamiltonians $\hat{H}/\hbar=\omega(\hat{f}_{1}^{\dagger}\hat{f}_{1}+\hat{f}_{2}^{\dagger}\hat{f}_{2})+|\kappa|(\hat{f}_{1}^{\dagger}\hat{f}_{2}+\hat{f}_{2}^{\dagger}\hat{f}_{1})/2$ and $\hat{H}/\hbar=\omega(\hat{b}_{1}^{\dagger}\hat{b}_{1}+\hat{b}_{2}^{\dagger}\hat{b}_{2})+|\kappa|(\hat{b}_{1}^{\dagger}\hat{b}_{2}+\hat{b}_{2}^{\dagger}\hat{b}_{1})/2$ for purely fermionic and bosonic systems in second quantization, respectively. The corresponding annihilation and creation operators satisfy the fundamental (anti-)commutation relations $\{\hat{f}_{j},\hat{f}_{j^{\prime}}\}=0=[\hat{b}_{j},\hat{b}_{j^{\prime}}]$ and $\{\hat{f}_{j},\hat{f}_{j^{\prime}}^{\dagger}\}=\delta_{j,j^{\prime}}=[\hat{b}_{j},\hat{b}_{j^{\prime}}^{\dagger}]$ for $j,j^{\prime}\in\{1,2\}$. Furthermore, we find a diagonalization of the Hamiltonians via the eigenmodes
$\hat{f}_{\pm}=\frac{\hat{f}_{1}\pm\hat{f}_{2}}{\sqrt{2}}\quad\text{and}\quad\hat{b}_{\pm}=\frac{\hat{b}_{1}\pm\hat{b}_{2}}{\sqrt{2}},$ (22)
which remain fermionic and bosonic, respectively.
For fermions, we find the Fock-basis expansion of $N$-particle states as
$|0,0\rangle_{\pm}={}$ $|0,0\rangle_{\mathrm{FF}}=|\mathrm{vac}\rangle,$ (23)
$|1,0\rangle_{\pm}={}$ $\frac{|1,0\rangle_{\mathrm{FF}}+|0,1\rangle_{\mathrm{FF}}}{\sqrt{2}},$
$|0,1\rangle_{\pm}={}$ $\frac{|1,0\rangle_{\mathrm{FF}}-|0,1\rangle_{\mathrm{FF}}}{\sqrt{2}},$
$|1,1\rangle_{\pm}={}$ $|1,1\rangle_{\mathrm{FF}},$
containing at most one excitation per mode because of the exclusion principle—likewise, the nilpotent operators $\hat{f}_{j}^{{\dagger}2}=0$. In addition, for bosons, we find the orthonormal $N$-particle states
$|m,n\rangle_{\pm}=\frac{\hat{b}_{+}^{{\dagger}m}\hat{b}_{-}^{{\dagger}n}}{\sqrt{m!n!}}|\mathrm{vac}\rangle=\sum_{s=0}^{N}q_{s|m,n}|s,N-s\rangle_{\mathrm{BB}},$ (24)
with
$q_{s|m,n}={}$ $\sqrt{\frac{m!n!s!(N-s)!}{2^{N}}}$ (25)
×$\times\sum_{l=\max\{0,s-m\}}^{\min\{s,n\}}\frac{(-1)^{n-l}}{l!(n-l)!(s-l)!(l+m-k)!}$
and $m,n\in\mathbb{N}$.
The vacuum states are in all cases identical. Also, the single-particle states, $(|0,1\rangle_{XY}\pm|1,0\rangle_{XY})/\sqrt{2}$, do not show differences for independent modes (i.e., $XY\in\{\mathrm{FF},\mathrm{BB}\}$) and dependent degrees of freedom, $XY=\mathrm{FB}$. Interesting differences appear for $N>1$. For instance, the two-particle states for two-mode fermions are always of a product form, regardless the mode decomposition. Thus, in the following, we continue our comparison of only quasi-particles and bosons for $N=2$.
Figure 3: Comparison of excitations of two bosons and two quasi-particles. The left plot shows the original space spanned by two bosons ($\mathrm{BB}$) and the subspace of the fermion-boson system ($\mathrm{FB}$). The center plot shows the effect of the rotation to the superposition modes, resulting in tensor-product basis states in the purely bosonic case. The right plot shows the projection $\hat{P}$, while preserving normalization, to the constrained quasi-particle states. In the middle plot, all states on the sphere but $|2-n\rangle_{+}\otimes|n\rangle_{-}$ and $-(|2-n\rangle_{+}\otimes|n\rangle_{-})$ for $n\in\{0,1,2\}$ exhibit entanglement of independent degrees of freedom. However, for dependent degrees of freedom, every state on the circle except for the projected points $|2-n,n\rangle_{\mathrm{\pm}}$, together with their reflections, are the entangled ones in the right plot.
IV.2 Two bosons, quasi-particles, and physical projections
The purely bosonic case yields the $\pm$-product states, with the explicit expansion
$|2\rangle_{+}\otimes|0\rangle_{-}={}$ $\frac{1}{2}\left(|2,0\rangle_{\mathrm{BB}}+\sqrt{2}|2,0\rangle_{\mathrm{BB}}+|0,2\rangle_{\mathrm{BB}}\right),$ (26)
$|1\rangle_{+}\otimes|1\rangle_{-}={}$ $\frac{1}{\sqrt{2}}\left(|2,0\rangle_{\mathrm{BB}}-|0,2\rangle_{\mathrm{BB}}\right),$
$|0\rangle_{+}\otimes|2\rangle_{-}={}$ $\frac{1}{2}\left(|2,0\rangle_{\mathrm{BB}}-\sqrt{2}|2,0\rangle_{\mathrm{BB}}+|0,2\rangle_{\mathrm{BB}}\right).$
For comparison, the explicit expansion of all pure factorizable two-quasi-particle states reads
$|2,0\rangle_{\pm}={}$ $\frac{1}{\sqrt{3}}\left(\sqrt{2}|1,1\rangle_{\mathrm{FB}}+|0,2\rangle_{\mathrm{FB}}\right),$ (27)
$|1,1\rangle_{\pm}={}$ $\left(-|0,2\rangle_{\mathrm{FB}}\right),$
$|0,2\rangle_{\pm}={}$ $\frac{1}{\sqrt{3}}\left(-\sqrt{2}|1,1\rangle_{\mathrm{FB}}+|0,2\rangle_{\mathrm{FB}}\right).$
Up to a normalization, all instances of both cases are only distinguished through the presence ($\mathrm{BB}$ case) and absence ($\mathrm{FB}$ case) of the $|2,0\rangle_{XY}$ component in the original mode decomposition. This makes sense as, for a single fermionic mode, we can identify
$\hat{f}=\hat{P}\hat{b}_{1}\hat{P}\quad\text{and}\quad\hat{b}=\hat{b}_{2},$ (28)
with a projection operator in Fock basis as
$\hat{P}=\sum_{n_{1}\in\{0,1\}}\sum_{n_{2}\in\mathbb{N}}|n_{1},n_{2}\rangle_{\mathrm{BB\,\,BB}}\langle n_{1},n_{2}|.$ (29)
One can straightforwardly check that a restriction of a bosonic annihilation operator to at most one excitation satisfies the properties of a fermionic annihilation operator. In other words, the above projection yields
$|m,n\rangle_{\pm}\propto\hat{P}|m\rangle_{\mathrm{+}}\otimes|n\rangle_{-}.$ (30)
Note that this relation holds true for $N=m+n>2$, too.
The cases of independent degrees of freedom (two bosonic modes) and dependent degrees of freedom (quasi-particle modes) are related via a projection. This projection also explains the linear dependence since $\hat{P}$ is not bijective and maps an orthonormal tensor-product basis $|m\rangle_{+}\otimes|n\rangle_{-}$ to a linearly dependent set of vectors $|m,n\rangle_{\pm}$. In general, the definition of non-entangled states in Eq. (11) now takes the simpler form
$|x,y\rangle_{\pm}=\hat{P}\,|x\rangle_{+}\otimes|y\rangle_{-},$ (31)
while choosing the scaling of the input states such that the projected result is properly normalized. In this form, the relation of non-entangled states $|x,y\rangle_{\pm}$ to tensor-product states, Eq. (31), becomes clearer. Also, the physical restriction of the composite system is achieved via the projection operator $\hat{P}$, projection onto the fermionic mode $\hat{f}$ as a part of both quasi-particle modes $\hat{p}_{\pm}$. See Fig. 3.
V Generalizations
While we solely considered a specific system of interest so far, we generalize the approach to entanglement between more general dependent degrees of freedom. This includes the treatment of multimode fermion-boson interfaces, a projection-based formalization for other systems, and the mentioning of another application in angular-momentum degrees of freedom.
V.1 Multimode generalization
The framework discussed thus far was based on one fermionic and one bosonic mode. Beyond that, suppose we have $\hat{f}_{j}$, with $j\in\{1,\ldots,M\}$, and $\hat{b}_{j}$ as annihilation operators of our quantized modes in an $2M$-mode composite system. These operators obey the anti-commutator and commutator relations,
$\{\hat{f}_{j},\hat{f}_{j^{\prime}}\}=0,$ $[\hat{b}_{j},\hat{b}_{j^{\prime}}]=0,$ (32)
$\{\hat{f}_{j},\hat{f}_{j^{\prime}}^{\dagger}\}=\delta_{j,j^{\prime}},\quad\text{and}$ $\quad[\hat{b}_{j},\hat{b}_{j^{\prime}}^{\dagger}]=\delta_{j,j^{\prime}}.$
A linear mode transformation reads
$\begin{bmatrix}\hat{p}_{+,1}\\ \vdots\\ \hat{p}_{+,M}\\ \hat{p}_{-,1}\\ \vdots\\ \hat{p}_{-,M}\end{bmatrix}=\begin{bmatrix}U_{+,F}&U_{+,B}\\ U_{-,F}&U_{-,B}\end{bmatrix}\begin{bmatrix}\hat{f}_{1}\\ \vdots\\ \hat{f}_{M}\\ \hat{b}_{1}\\ \vdots\\ \hat{b}_{M}\end{bmatrix},$ (33)
with matrices $U_{+,F},U_{+,B},U_{-,F},U_{-,B}\in\mathbb{C}^{M\times M}$.
Non-entangled states, akin to Eq. (11), now are defined as
$|x^{(1)},\ldots,x^{(2M)}\rangle_{\pm}=\prod_{\begin{subarray}{c}s\in\{+,-\},\\ j\in\{1,\ldots,M\}\end{subarray}}\left(\sum_{m\in\mathbb{N}}\tilde{x}_{m}^{(s,j)}\hat{p}_{s,j}^{{\dagger}m}\right)|\mathrm{vac}\rangle.$ (34)
A convex hull construction then allows us to express mixed non-entangled (i.e., separable) states as statistical mixtures
[…]
×$\times|x^{(1)},\ldots,x^{(2M)}\rangle_{\pm\,\,\pm}\langle x^{(1)},\ldots,x^{(2M)}|,$
where $P$ denotes a joint probability distribution to account for classical correlations. Whenever a state $\hat{\varrho}$ is not in a separable configuration, $\hat{\varrho}\neq\hat{\rho}_{\mathrm{sep.}}$, we have an inseparable state that exhibits quantum correlations.
Formally, the above definition of multimode non-entangled states for dependent degrees of freedom is identical to that for independent degrees of freedom. Hence, it is consistent with the existing approaches. However, as we have shown in Sec. III for $M=2$, the entanglement and disentanglement of states can deviate from the common expectation.
We also remark on the fact that the mapping in Eq. (33) does not have to be a unitary one. For instance, this allows us to describe product and entangled states in partially overlapping modes, e.g., $\hat{p}_{-,1}=\hat{b}_{1}$ and $\hat{p}_{-,2}=(\hat{b}_{1}+\hat{b}_{2})/\sqrt{2}$, which is interesting in the context of partial coherence and entanglement in quantum optics [37, 38].
V.2 Projected product states
As shown in this work, the entanglement characteristics of quasi-particles with dependent degrees of freedom can be determined through non-entangling states that are projected product states, Eq. (30). We can formalize this approach. Let $\mathcal{H}=\hat{P}(\mathcal{H}_{X}\otimes\mathcal{H}_{Y})$ be a composite Hilbert space subject to a constraint given by the projection $\hat{P}$, with $\hat{P}^{2}=\hat{P}=\hat{P}^{\dagger}$. A pure non-entangled state can be defined as
$|x,y\rangle_{XY}=\hat{P}\left(|x\rangle_{X}\otimes|y\rangle_{Y}\right),$ (36)
where $|x\rangle_{X}\in\mathcal{H}_{X}$ and $|y\rangle_{Y}\in\mathcal{H}_{Y}$ are local states of the joint system. Note that this definition is identical with the one for a factorizable state for independent degrees of freedom, where $\hat{P}=\hat{1}$. An entangled state $|\Psi\rangle$ does not admit the decomposition from Eq. (36), $|\Psi\rangle\neq\hat{P}\,(|x\rangle_{X}\otimes|y\rangle_{Y})$.
Furthermore, a mixed, classically correlated state reads
[…]
And a state that does not obey this form is consequently inseparable, i.e., entangled. Also note that this definition straightforwardly generalizes to multipartite systems. Therefore, a separable state is defined as a state that obeys the physical constraint (given by the projector $\hat{P}$) and only contains classical correlations (via the joint probability distribution $P$) of local tensor-product states $|x\rangle_{X}\otimes|y\rangle_{Y}$.
In addition to the quasi-particle description, whose projection can be found in Eq. (29), another example is in order. Consider the degree of freedom pertaining to the total angular momentum, $J^{2}$, and one for the angular momentum in $z$ direction, $J_{z}$. It is well known that the corresponding angular-momentum operators can be jointly diagonalized,
$\hat{J}^{2}|j,m\rangle_{J^{2}J_{z}}={}$ $\hbar^{2}j(j+1)|j,m\rangle_{J^{2}J_{z}}$ (38)
$\text{and}\quad\hat{J}_{z}|j,m\rangle_{J^{2}J_{z}}={}$ $\hbar m|j,m\rangle_{J^{2}J_{z}},$
$\text{with}\quad-j\leq m\leq j$ $\quad\text{and}\quad j\geq 0.$
From the latter restriction of values, the imposed physical constraint can be reformulated through the projection
$\hat{P}=\sum_{j:j\geq 0}|j\rangle_{J^{2}\,J^{2}}\langle j|\otimes\sum_{m:-j\leq m\leq j}|m\rangle_{J_{z}\,J_{z}}\langle m|,$ (39)
where $j$ and $m$ could, in principle, take arbitrary (half) integer values without the restrictions. The projection $\hat{P}$ then makes sure that the angular momentum in one direction never exceeds the total angular momentum, $|m|\ngtr j$. The above definition further allows us to define non-entangled states,
$|x,y\rangle_{J^{2}\,J_{z}}=\hat{P}\left(\sum_{j}x_{j}|j\rangle_{J^{2}}\right)\otimes\left(\sum_{m}y_{m}|j\rangle_{J_{z}}\right).$ (40)
For instance, the state $(|0,0\rangle_{J^{2}\,J_{z}}+|1,-1\rangle_{J^{2}\,J_{z}}+|1,0\rangle_{J^{2}\,J_{z}}+|1,1\rangle_{J^{2}\,J_{z}})/2$ is not of the form $|x\rangle_{J^{2}}\otimes|y\rangle_{J_{z}}$ itself, but it can be obtained via a projection of the local states $|x\rangle_{J^{2}}\propto|0\rangle_{J^{2}}+|1\rangle_{J^{2}}$ and $|y\rangle_{J_{z}}\propto|-1\rangle_{J_{z}}+|0\rangle_{J_{z}}+|-1\rangle_{J_{z}}$, in which the projection removes the unphysical parts $|0\rangle_{J^{2}}\otimes|\pm 1\rangle_{J_{z}}$. See also Ref. [39] for the relation to other notions of nonclassicality.
As a finial remark, it is worth mentioning that $\hat{P}$ does not have to be a projection operator for our above general definitions of separability for dependent degrees of freedom to work. Analogously to the aforementioned scenario of partially overlapping modes, we can generalize $\hat{P}$ to arbitrary linear operators if required by a physical constraint that yields the composite system, $\mathcal{H}=\hat{P}(\mathcal{H}_{X}\otimes\mathcal{H}_{Y})$.
VI Conclusion
In this work, we explored the impact of dependent degrees of freedom on quantum entanglement properties. Defying common expectations, we showed that certain states are non-entangled in the case of dependent degrees of freedom which are entangled for independent notions of entanglement, and vice versa.
The main physical system to which we applied our approach were quasi-particle descriptions for light-matter quantum interfaces via a Jaynes–Cummings model. We showed that non-entangled quasi-particle states can be expressed through projections—accounting for physical constraints—that act on tensor-product states. The projection leads to linear dependencies, enabling us, for example, to construct NOON-type states which are not entangled, unlike their counterparts for independent degrees of freedom. Conversely, the true statement that eigenvectors of a local (and non-degenerate) Hamiltonian are necessarily factorizable is generally not true for dependent degrees of freedom.
Furthermore, we compared our fermion-boson-composite quasi-particle description with purely bosonic and fermionic systems. For a single excitation, no distinction could be made between the different cases. However, we were able to identify unique entanglement characteristics for the different physical scenarios when considering at least two (quasi-)particles. Eventually, a generalization to multipartite entanglement of dependent degrees of freedom, including the treatment of mixed states and physical systems other than quasi-particles, were established.
Therefore, a framework was put forward to characterize quantum correlations between dependent degrees of freedom. This allows us to go beyond, for instance, the assumed well-separated nature of sender and receiver in quantum communication to accommodate for physical constraints not contributing to a quantum protocol as a quantum resource.
Acknowledgements.
The authors acknowledge funding through the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) via the transregional collaborative research center TRR 142 (Projects A04 and C10, Grant No. 231447078). This work was further supported through the Ministerium für Kultur und Wissenschaft des Landes Nordrhein-Westfalen through the project PhoQC: Photonisches Quantencomputing.